OpenStudy (anonymous):

140 mi in 3.5 h a. 40 mi per h c. 0.03 mi per h b. 143.5 mi per h d. 35 mi per h

OpenStudy (anonymous):

@nirmalnema

OpenStudy (nirmalnema):

@xxkerstie.bbyxx do you need to fine the speed??

OpenStudy (nirmalnema):

mi...is what..??

OpenStudy (anonymous):

find each unit rate

OpenStudy (texaschic101):

3.5/140 = 1/x cross multiply 3.5x = 140 x = 140/3.5 x = 40 miles per hr.

OpenStudy (nirmalnema):

the ans is c /....that is 40 min per hr.... 140mi in 3.5 hr so in 1 hr 140/3.5=40 mi per hr